OpenStudy (xmissalycatx):

Just a quick question to refresh my memory on exponents.. @mathmale @kainui

OpenStudy (xmissalycatx):

If you have something like \[3x ^{-4}\] is it still written like this, or would you do something to keep it from being a negative exponent?

OpenStudy (mathmale):

Your \[3x ^{-4}\]

OpenStudy (mathmale):

is fine as is. However, if you are told that your exponents must be positive, use the following rule to rewrite that expression:\[x ^{-a}=\frac{ 1 }{ x^a }\]

OpenStudy (mathmale):

Therefore, \[x ^{-4}=?\]

OpenStudy (xmissalycatx):

So I could write it as \[\frac{ 3 }{ x^4 }\]

OpenStudy (xmissalycatx):

Because that is what I did originally.

OpenStudy (mathmale):

Yes, your original \[3x ^{-4}=\frac{ 3 }{ x^4 }\]

OpenStudy (xmissalycatx):

Okay, thank you so much! :)

OpenStudy (mathmale):

;)